OpenStudy (anonymous):

Please help me solve this. I am not looking for the answer because I don't want to cheat, I promise. What is: 9c + 4 = -23?

OpenStudy (anonymous):

ok so to start off you want the variable, c, on its own, you would start by moving the 4 to the other side, you do this by subtracting 4 from both sides 9c+4-4= -23-4

OpenStudy (anonymous):

Okay, so it would be 9c + 0 = 19?

OpenStudy (anonymous):

close but not quite, your right about 9c+0 but if it is a negative 23 you are going further negative by subtracting 4

OpenStudy (anonymous):

I'm sorry, but I'm still confused! I know, I'm a pain, haha.

OpenStudy (anonymous):

dont worry about it, my teacher gave us an illistration, think of the negative numbers as bruises and the positive as bandages, so if you had 5 bandages and 4 bruises you would have 1 bandage 5-4=1, but if you had 5 bruises and then 4 more bruises you would have 9 bruises -5-4=-9 using this if you had 23 bruises (-23) and then 4 more bruises (-4) how many bruises would you have? -23-4=?

OpenStudy (anonymous):

Oh... okay, so it would be -27?

OpenStudy (anonymous):

yeap!

OpenStudy (anonymous):

So, 9c+4-4=-23-4 is... 9c+0=-27...

OpenStudy (anonymous):

exactly

OpenStudy (anonymous):

the next thing you need to do is bring the 9 to the other side, in order to bring something to the other side you need to do the opposite of it. because the equation is c times 9 you need to divide by 9 to get rid of it \[9c \div 9=-27\div9\]

OpenStudy (anonymous):

So it's 1c = -3?

OpenStudy (anonymous):

Or c = -3?

OpenStudy (anonymous):

you got it!

OpenStudy (anonymous):

Thank you so much! I am not good at math and the course that I am taking doesn't explain anything very well. :D

OpenStudy (anonymous):

no problem! ill be glad to help anytime i can

OpenStudy (anonymous):

Thanks!